write the equation of a line parallel to the given line but passing through the given point. Show work
y=-x+3;(-2,-2)

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

y = -1 x + 3
Divide by 1
y = -1 x + 3
Compare this equation with y=mx+b
slope m = -1
The slope of a line parallel to the above line will be the same
The slope of the required line will be -1
m= -1 ,point ( -2 , -2 )
Find b by plugging the values of m & the point in
y=mx+b
-2 = 2 + b
b= -4
m= -1
Plug value of the slope and b in y = mx +b
The required equation is y = -1 x -4

m.ananth@hotmail.ca

Answer by stanbon(75887)

write the equation of a line parallel to the given line but passing through the given point. Show work
y=-x+3 ; (-2,-2)
-------
The given equation has slope = -1
So every line parallel to it has slope = -1
-------
Form: y = mx + b
Solve for "b":
-2 = -1*-2 + b
b = -4
---------
Equation:
y = -x -4
==================
Cheers,
Stan H.
==================